?15c) If the percentage of all male births is 51%, what is the probablility that a five child family will have at least one boy?
Answer #1

P(male birth) = 0.51

P(female birth) = 1 - 0.51 = 0.49

P(at least one boy) = 1 - P(all are girl)

                               = 1 - 0.495

                               = 1 - 0.0282

                               = 0.9718

                               = 97.18%